OpenStudy (anonymous):

simplify tan^2(x)-tan^2(x)sin^2(x)

ganeshie8 (ganeshie8):

start by factoring out `tan^2(x)`

ganeshie8 (ganeshie8):

tan^2(x)-tan^2(x)sin^2(x) tan^2(x)[1 - sin^2(x)]

OpenStudy (anonymous):

what trig identity allows one to factor out the tan^2x? , I'm having trouble grasping the concept

OpenStudy (zzr0ck3r):

That is not an identity. Its just the distribution property a(1-b) = a-ab for you a = tan^2(x) and b = sin^2(x)

OpenStudy (anonymous):

ah okay,

OpenStudy (ranga):

Use trig identity to simplify what ganeshie did before: 1 - sin^2(x) = ? Also write tan^2(x) as sin^2(x) / cos^2(x). Then simplify.

OpenStudy (anonymous):

got it. thanks!

OpenStudy (ranga):

You are welcome.